# UGround: Towards Unified Visual Grounding with Unrolled Transformers

### v1

@@ 1 Introduction
Visual grounding aims to align referring expressions with their corresponding regions in an image or video  (Mao et al., 2016; Yu et al., 2016). Despite existing efforts  (Zou et al., 2023a; Liu et al., 2024b; Zou et al., 2023b; Liang et al., 2023; Liu et al., 2023; Wang et al., 2022), it remains largely unexplored under the newly emerging task settings  (Lai et al., 2024; Wu et al., 2024; Xia et al., 2024; Rasheed et al., 2024; Ren et al., 2024; Qian et al., 2025). As shown in Table 1, the shifts across recent tasks settings reveal how visual grounding has evolved, from referring to objects by an explicit mention (RES (Rasheed et al., 2024; Ren et al., 2024)) to reasoning over descriptive language in an implicit fashion (RS (Lai et al., 2024; Yang et al., 2023)); from processing single-target queries per instance to handling multi-target scenarios (gRES (Xia et al., 2024), Multi-RS (Ren et al., 2024)); and from responding solely to positive queries to rejecting false premises with empty targets (FP-RES (Wu et al., 2024; Xia et al., 2024; Qian et al., 2025)). Such shifts reflect the intrinsic properties of tasks, which we term “attribute variation”.
However, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, diverging from prior works (Ren et al., 2024; Rasheed et al., 2024; Wei et al., 2025) that primarily focus on versatile capabilities. Generally, perception systems engaging with real-world scenarios are expected to reason over implicit instructions (cognition), interact with one or multiple targets (generality), and appropriately reject queries when necessary (safety). Built upon attribute basis described above, versatility-level extensions, such as grounding in videos (Wei et al., 2025) or conversations (Rasheed et al., 2024), find their footing. This leads us to ask: Can we design a unified architecture to bridge this gap?
Figure 1: Prior works typically use <SEG> token embeddings from the last hidden layer as the prompt in (a). In contrast, we leverage the similarity map in Eq. (6), generated from <SEG> and image token embeddings across dynamically selected transformer layers, as the prompt in (b). Dynamic layer selection allows the similarity map to slide across transformer layers in (c). Due to the sequential nature of transformers, $\ell_{L}-\ell_{k}$ layers are skipped, enabling a direct connection with SAM in a single forward step $\mathcal{T}_{t}$ (“skip-connection-like”). Across multiple forward steps $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, the connectivity to SAM varies dynamically, with only one path activated at each step (“dropout-like”).
@@
To this end, we present UGround, a Unified visual Grounding paradigm that dynamically selects intermediate layers across Unrolled transformers as “mask as prompt”. Central to our UGround is Policy-Prompted Masking (PPM), which consists of two key components: Stochastic Skip Connections (SSC) and Mask as Prompt (MasP). As shown in Fig. 1(c), SSC is a stochastic reinforcement learning policy that allows each <SEG> token to slide across unrolled transformer layers, reinforcing dynamic layer selection at which it connects to the vision model (e.g., SAM) in a skip-connection fashion. Remarkably, SSC enables intermediate layers to skip subsequent ones and connect directly with SAM within each forward $\mathcal{T}_{t}$ (“skip-connection-like”). Across forward $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, such mechanism virtually connects all layers to SAM while activating only one pass at a time (“dropout-like”). Given the selected hidden layer, MasP uses the similarity map derived from the <SEG> token and image tokens as a soft logit mask to prompt SAM for mask generation, offering explicit spatial cues through its activation regions in the forward. Particularly, the similarity map is differentiable, beyond its implicit optimization via SAM backpropagation, we further impose a cross-entropy loss against the ground-truth mask to explicitly guide spatial cues in the backward. In summary, our contributions are threefold:
@@
  • We have “unrolled” stacked transformers, letting vision decoder tap into intermediate layers in a stochastic skip-connection fashion, in the way of “mask as prompt”. Importantly, we explicitly supervise the prompted mask (similarity map) against the ground truth mask to further guide the model on where to attend.
  • We conduct extensive experiments on the recently proposed ReasonSeg dataset, the classic RefCOCO(+/g) dataset, and gRefCOCO(+/g). Our UGround outperforms the state of the art, achieving up to 9.0% gains in cIoU on ReasonSeg and 12.1% gains in N-acc on gRefCOCO(+/g) val sets.
@@ 2 Related Work
Large Multimodal Models. Depending on the trade-offs between computational efficiency and performance ceiling, we categorize LMMs into two groups: (1) Fusion-centric models, which aim for deeply fine-grained multimodal interaction by projecting visual, textual, and other modalities into a shared high-dimensional space with cross-attention mechanisms (Alayrac et al., 2022; Wang et al., 2023; Pi et al., 2023; Lv et al., 2023; Zhang et al., 2024a). Flamingo (Alayrac et al., 2022) unlocks few-shot visual reasoning by bridging a frozen language model and a vision encoder with gated cross-attention. VisionLLM (Wang et al., 2023) conditions a language model on image features via cross-attention to handle diverse vision-centric tasks, generating task-specific tokens for open-ended reasoning. GPT4RoI (Zhang et al., 2024a) enables region-of-interest (RoI) reasoning through RoI-aware cross‑attention, allowing users to reference any image region in dialogue. Kosmos-2 (Peng et al., 2024) and DetGPT (Pi et al., 2023) equip LMMs with grounding capabilities. (2) Alignment-centric models, which employ a divide-and-conquer strategy, projecting the outputs of frozen or lightly fine-tuned unimodal encoders (e.g., ViT for vision, LLaMA/GPT for text) into a shared latent space via a lightweight alignment adapter (e.g., projectors, Q-Formers) (Li et al., 2023; Ye et al., 2023; Li et al., 2025; Liu et al., 2024a; Zhu et al., 2024). Otter (Li et al., 2025) instruction-tunes a Flamingo-style vision–language model on the MIMIC-IT dataset. LLaVA (Liu et al., 2024a) and MiniGPT-4 (Zhu et al., 2024) align a pretrained vision encoder with a frozen LLM using lightweight adapters for open-ended, instruction-following visual reasoning. Unlike LLaVA and MiniGPT-4, which rely on linear projections, BLIP-2 (Li et al., 2023) pioneers the Q-Former to distill visual features into compact queries for alignment with a frozen LLM. mPLUG-OWL (Ye et al., 2023) utilizes a visual abstractor to summarize global image features into condensed semantic tokens. In contrast, our UGround builds upon LLaVA (Liu et al., 2024a) for visual autoregressive modeling.
@@
Unified Segmentation Models. Depending on the semantic granularity of joint tasks, we divide literature into two groups: Versatility-oriented segmentation, which aims to unify multiple functional properties within a single model (Wei et al., 2025; Cheng et al., 2022; Li et al., 2024; Zhang et al., 2024b; Bai et al., 2024; Rasheed et al., 2024), as opposed to prior works dedicated exclusively to semantic (Long & Shelhamer, 2015; Ronneberger et al., 2015; Badrinarayanan et al., 2017), instance (He et al., 2017; Liu et al., 2018; Carion et al., 2020), or panoptic segmentation (Kirillov et al., 2019b; Xiong et al., 2019; Kirillov et al., 2019a). Mask2Former (Cheng et al., 2022) pioneers a unified approach for image-level semantic, instance, and panoptic segmentation. OMG-Seg (Li et al., 2024) extends Mask2Former to their video-level counterparts. HyperSeg (Wei et al., 2025) and OMG-LLAVA (Zhang et al., 2024b) distinguish themselves from previous purely visual models (Cheng et al., 2022; Li et al., 2024) by leveraging LMMs to generate interactive text. VideoLISA (Bai et al., 2024) and HyperSeg (Wei et al., 2025) inject complex reasoning capabilities into generic video segmentation models. (2) Attribute-oriented segmentation, which seeks to unify the intrinsic property variations within the task itself, spanning from explicit expressions to implicit instructions, one target to many targets, and positive queries to false premises. LISA (Lai et al., 2024) elevates RES to RS by leveraging self-reasoning capabilities of LMMs. SESAME (Wu et al., 2024) and READ (Qian et al., 2025) extend LISA by rejecting false premises. GSVA (Xia et al., 2024) and PixelLM (Ren et al., 2024) generalize RES to gRES by supporting multiple targets. Closest to our work, PixelLM (Ren et al., 2024) cannot tackle empty targets, while GSVA (Xia et al., 2024) fails to handle multi-target reasoning. Attribute-oriented unification is more fundamental, as it can be applied from a versatile perspective to any specialized task. Yet, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, we therefore present UGround.
@@ 3.1 Revisiting
(a) cloU on ReasonSeg val set across layers 1-40.
(b) Gauss dice loss on ReasonSeg val set across layers 1-40.
Figure 2: (2(a)) shows that the dynamic layer selection strategy outperforms the fixed last hidden layer strategy across all middle layers. (2(b)) plots the loss of similarity maps against the soft ground-truth mask, demonstrating faster convergence at intermediate layers.
@@ 3.2 Analysis
Why Dynamic Layer Selection. To analyze the impact of intermediate layers, we train UGround on the ReasonSeg train set with two strategies: (1) relying solely on the last hidden layer ($\ell=L$), and (2) dynamically selecting from the intermediate layers ($\ell\in\{1,\dots,L-1\}$). Qualitatively, we visualize the similarity maps across layers 1–40 of $\mathcal{G}_{\mathcal{T}}$ in Fig. 6 (Appendix 5). We observe that training with the fixed last hidden layer yields noisy similarity maps in intermediate layers, which eventually undermine the discriminative capacity of the final abstracted representations (6(a)). In contrast, with dynamic layer selection, the middle layers can directly align with the ground-truth mask, yielding a more discriminative foreground (6(b)). Quantitatively, we plot the performance curves of the embeddings across layers 1–40, which reveal that each layer from 10–40 outperforms the fixed last hidden layer strategy (2(a)). We also plot the per-layer similarity map loss relative to the ground-truth mask, which indicates that dynamic layer selection accelerates training convergence (2(b)).
Table 2: Quantitative analysis of the similarity maps on the ReasonSeg test split. $\mathcal{M}$prompt indicates using the mask as a prompt for the original SAM (Kirillov et al., 2023), while $\mathcal{M}$IoU quantifies the overlap (IoU) between the similarity map and the ground-truth mask. <SEG>prompt refers to using the <SEG> token as prompt for the adapted SAM, i.e., LISA (Lai et al., 2024).
Method $\mathcal{M}$prompt $\mathcal{S}$IoU <SEG>prompt
@@
Why Similarity as Mask. To quantitatively assess the spatial cues encoded in the similarity maps, we use them as logit masks to prompt the original SAM (Kirillov et al., 2023) for mask generation. Table 2 reveals that $\mathcal{M}$prompt is potentially capable of generating segmentation masks solely from similarity maps ($17.0\%$ vs. $48.4\%$). To further quantify the alignment between the activated regions of the similarity map shown in Fig. 6(a) and the ground-truth mask in Fig. 3, we employ the grid search-based Intersection over Union (IoU) from READ (Qian et al., 2025) to evaluate its consistency. In Table 2, $\mathcal{M}$IoU indicates that activations in the similarity map align closely with the ground-truth mask, as evidenced by a 4.3% higher cIoU on the ReasonSeg test set ($35.0\%\ \text{vs.}\ 30.7\%$) in the $3^{\text{rd}}$ row.
Summary. Analysis of the similarity maps reveals the following: (a) Similarity maps in LMMs are highly consistent with the ground-truth mask, suggesting that they can potentially serve as prompts for generating the segmentation mask. (b) Dynamic layer selection allows the intermediate layers to directly interact with SAM, resulting in a more discriminative representation, as reflected by the activated regions in the similarity maps. Motivated by the insights from the similarity maps, we leverage it in three ways: as a logit mask to prompt SAM (prompt), as a loss constraint to guide the model on where to “attend” (constraint), and as a policy reward for dynamic layer selection (signal) (see Appendix 5 for further analysis).
@@ 4.1 Policy-Prompted Masking
$\pi_{\theta}(\ell\mid\mathcal{H}_{t^{*}})=\frac{\exp(s_{\ell})}{\sum_{j=1}^{L}\exp(s_{j})},s_{\ell}=\bm{h}_{t^{*}}^{(\ell)}\cdot\mathbf{w}_{\ell},$ (4)
@@
where the log-probability of the sampled action $a\equiv\ell^{*}$ is weighted by the advantage rbr-b.
To get the reward $r$, we first compute the similarity map $\mathcal{M}$ between the <SEG> token and image tokens at layer $\ell$ and its corresponding soft mask $\mathcal{M}_{\sigma}$. Consider a sequence of discrete image tokens $\mathbf{z}=(z_{1},z_{2},\dots,z_{k})$ in the input sequence $\mathbf{x}$, where $k$ is the number of image tokens. The image token embeddings across layers are then denoted as $\mathcal{H}_{\mathbf{z}}=\{\bm{h}_{z}^{(\ell)}\mid z\in\mathbf{z},\ \ell=1,\dots,L\}\in\mathbb{R}^{k\times L\times d}$, where $\bm{h}_{z}^{(\ell)}$ is the embedding of the $z$-th image token at layer $\ell$. Given the <SEG> token embedding $\bm{h}_{t^{*}}^{(\ell)}\in\mathbb{R}^{d}$ and the image token embeddings $\mathcal{H}_{\mathbf{z}}^{(\ell)}\in\mathbb{R}^{k\times d}$ at the selected layer $\ell$, the similarity score can be computed as
@@
where each element $\mathcal{S}_{i}^{(\ell)}=(h_{z_{i}}^{(\ell)})^{\top}h_{t^{*}}^{(\ell)}$ represents the similarity between the <SEG> token and the $i$-th selected image token at layer $\ell$. $\mathcal{S}^{(\ell)}$ are arranged in a 2D grid, and interpolated to obtain the soft logit mask $\mathcal{M}\in[0,1]^{H\times W}$. Let the ground-truth mask be $M\in\{0,1\}^{H\times W}$, which we view as a function defined over the continuous 2D plane $\mathbb{R}^{2}$. By applying Gaussian smoothing, we obtain a continuous-valued function $M_{\sigma}:\mathbb{R}^{2}\to[0,1]$, where $M_{\sigma}$ represents the soft label or heatmap. Formally, $M_{\sigma}$ can be expressed as a convolution integral as $M_{\sigma}(u,v)=\iint_{\mathbb{R}^{2}}M(x,y)\,G_{\sigma}(u-x,v-y)\,dx\,dy,$ where $G_{\sigma}(x,y)\propto\exp(-\frac{x^{2}+y^{2}}{2\sigma^{2}})$, is a 2D Gaussian kernel with standard deviation $\sigma$. The reward for REINFORCE is computed using the similarity map $\mathcal{M}$ and the soft label $M_{\sigma}$ as $r=-\big(\mathcal{L}_{bce}(\mathcal{M},M_{\sigma})+\mathcal{L}_{dice}(\mathcal{M},M_{\sigma})\big),$ where $\mathcal{L}_{bce}$ and $\mathcal{L}_{dice}$ denote the Gaussian BCE and Dice losses defined over the similarity map $\mathcal{M}$. The policy loss is then defined as

### v2

@@ 1 Introduction
Visual grounding aims to align referring expressions with their corresponding regions in an image or video  (Mao et al., 2016; Yu et al., 2016). Despite existing efforts  (Zou et al., 2023a; Liu et al., 2024b; Zou et al., 2023b; Liang et al., 2023; Liu et al., 2023; Wang et al., 2022), it remains largely unexplored under the newly emerging task settings  (Lai et al., 2024; Wu et al., 2024; Xia et al., 2024; Rasheed et al., 2024; Ren et al., 2024; Qian et al., 2025). As shown in Table 1, the shifts across recent tasks settings reveal how visual grounding has evolved, from referring to objects by an explicit mention (RES (Rasheed et al., 2024; Ren et al., 2024)) to reasoning over descriptive language in an implicit fashion (RS (Lai et al., 2024; Yang et al., 2023)); from processing single-target queries per instance to handling multi-target scenarios (gRES (Xia et al., 2024), Multi-RS (Ren et al., 2024)); and from responding solely to positive queries to rejecting false premises with empty targets (FP-RES (Wu et al., 2024; Xia et al., 2024)). Such shifts reflect the intrinsic properties of tasks, which we term “attribute variation”.
@@
However, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, diverging from prior works (Ren et al., 2024; Rasheed et al., 2024) that primarily focus on versatile capabilities. Generally, perception systems engaging with real-world scenarios are expected to reason over implicit instructions (cognition), interact with one or multiple targets (generality), and appropriately reject queries when necessary (safety). Built upon attribute basis described above, versatility-level extensions, such as grounding in videos (Wei et al., 2025) or conversations (Rasheed et al., 2024), find their footing. This leads us to ask: Can we design a unified architecture to bridge this gap?
Figure 1: Prior works typically use <SEG> token embeddings from the last hidden layer as the prompt in (a). In contrast, we leverage the similarity map in Eq. (6), generated from <SEG> and image token embeddings across dynamically selected transformer layers, as the prompt in (b). Dynamic layer selection allows the similarity map to slide across transformer layers in (c). Due to the sequential nature of transformers, $\ell_{L}-\ell_{k}$ layers are skipped, enabling a direct connection with SAM in a single forward step $\mathcal{T}_{t}$ (“skip-connection-like”). Across multiple forward steps $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, the connectivity to SAM varies , with only one path activated at each step (“dropout-like”).
@@
To this end, we present UGround, a Unified visual Grounding paradigm that dynamically selects intermediate layers across Unrolled transformers as “mask as prompt”. Central to our UGround is Policy-Prompted Masking (PPM), which consists of two key components: Stochastic Skip Connections (SSC) and Mask as Prompt (MasP). As shown in Fig. 1(c), SSC is a stochastic reinforcement learning policy that allows each <SEG> token to slide across unrolled transformer layers, reinforcing dynamic layer selection at which it connects to the vision model (e.g., SAM) in a skip-connection fashion. Remarkably, SSC enables intermediate layers to skip subsequent ones and connect directly with SAM within each forward $\mathcal{T}_{t}$ (“skip-connection-like”). Across forward $\mathcal{T}_{1}$ to $\mathcal{T}_{T}$, such mechanism virtually connects all layers to SAM while activating only one pass at a time (“dropout-like”). Given the selected hidden layer, MasP uses the similarity map derived from the <SEG> token and image tokens as a soft logit mask to prompt SAM for mask generation, offering explicit spatial cues through its activation regions in the forward. Particularly, the similarity map is differentiable, beyond its implicit optimization via SAM backpropagation, we further impose a cross-entropy loss against the ground-truth mask to explicitly guide spatial cues in the backward. Our contributions are threefold:
@@
  • We have “unrolled” stacked transformers, letting vision decoder tap into intermediate layers in a stochastic skip-connection fashion, in the way of “mask as prompt”. Importantly, we explicitly supervise the prompted mask (similarity map) against the ground-truth mask to further guide the model on where to attend.
  • We conduct extensive experiments on the recently proposed ReasonSeg, the classic RefCOCO(+/g), and gRefCOCO(+/g) dataset. UGround outperforms the state of the art, achieving up to 9.0% gains in cIoU on ReasonSeg and 12.1% gains in N-acc on gRefCOCO(+/g) val sets.
@@ 2 Related Work
Large Multimodal Models. Depending on the trade-offs between computational efficiency and performance ceiling, we categorize LMMs into two groups: (1) Fusion-centric models, which aim for deeply fine-grained multimodal interaction by projecting visual, textual, and other modalities into a shared high-dimensional space with cross-attention mechanisms (Alayrac et al., 2022; Wang et al., 2023; Pi et al., 2023; Lv et al., 2023; Zhang et al., 2024a). Flamingo (Alayrac et al., 2022) unlocks few-shot visual reasoning by bridging a frozen language model and a vision encoder with gated cross-attention. VisionLLM (Wang et al., 2023) conditions a language model on image features via cross-attention to handle diverse vision-centric tasks, generating task-specific tokens for open-ended reasoning. GPT4RoI (Zhang et al., 2024a) enables region-of-interest (RoI) reasoning through RoI-aware cross‑attention, allowing users to reference any image region in dialogue. Kosmos-2 (Peng et al., 2024) and DetGPT (Pi et al., 2023) equip LMMs with grounding capabilities. (2) Alignment-centric models, which employ a divide-and-conquer strategy, projecting the outputs of frozen or lightly fine-tuned unimodal encoders (e.g., ViT for vision, LLaMA/GPT for text) into a shared latent space via a lightweight alignment adapter (e.g., projectors, Q-Formers) (Li et al., 2023; Ye et al., 2023; Li et al., 2025; Liu et al., 2024a; Zhu et al., 2024). Otter (Li et al., 2025) instruction-tunes a Flamingo-style vision–language model on the MIMIC-IT dataset. LLaVA (Liu et al., 2024a) and MiniGPT-4 (Zhu et al., 2024) align a pretrained vision encoder with a frozen LLM using lightweight adapters for open-ended, instruction-following visual reasoning. Unlike LLaVA and MiniGPT-4, which rely on linear projections, BLIP-2 (Li et al., 2023) pioneers the Q-Former to distill visual features into compact queries for alignment with a frozen LLM. mPLUG-OWL (Ye et al., 2023) utilizes a visual abstractor to summarize global image features into condensed semantic tokens. In contrast, our UGround builds upon LLaVA (Liu et al., 2024a) for visual autoregressive modeling and reasoning.
@@
Unified Segmentation Models. Depending on the semantic granularity of joint tasks, we divide literature into two groups: Versatility-oriented segmentation, which aims to unify multiple functional properties within a single model (Wei et al., 2025; Cheng et al., 2022; Li et al., 2024; Zhang et al., 2024b; Bai et al., 2024; Rasheed et al., 2024), as opposed to prior works dedicated exclusively to semantic (Long and Shelhamer, 2015; Ronneberger et al., 2015; Badrinarayanan et al., 2017), instance (He et al., 2017; Liu et al., 2018; Carion et al., 2020), or panoptic segmentation (Kirillov et al., 2019b; Xiong et al., 2019; Kirillov et al., 2019a). Mask2Former (Cheng et al., 2022) pioneers a unified approach for image-level semantic, instance, and panoptic segmentation. OMG-Seg (Li et al., 2024) extends Mask2Former to their video-level counterparts. HyperSeg (Wei et al., 2025) and OMG-LLaVA (Zhang et al., 2024b) distinguish themselves from previous purely visual models (Cheng et al., 2022; Li et al., 2024) by leveraging LMMs to generate interactive text. VideoLISA (Bai et al., 2024) and HyperSeg (Wei et al., 2025) inject complex reasoning capabilities into generic video segmentation models. (2) Attribute-oriented segmentation, which seeks to unify the intrinsic property variations within the task itself, spanning from explicit expressions to implicit instructions, one target to many targets, and positive queries to false premises. LISA (Lai et al., 2024) elevates RES to RS by leveraging self-reasoning capabilities of LMMs. SESAME (Wu et al., 2024) and READ (Qian et al., 2025) extend LISA by rejecting false premises. GSVA (Xia et al., 2024) and PixelLM (Ren et al., 2024) generalize RES to gRES by supporting multiple targets. Closest to our work, PixelLM (Ren et al., 2024) cannot tackle empty targets, while GSVA (Xia et al., 2024) fails to handle multi-target reasoning. Attribute-oriented unification is more fundamental, as it can be applied from a versatile perspective to any specialized task. Yet, we observe that little research has looked into unifying existing tasks within a single framework from an attribute perspective, we therefore present UGround.
@@ 3.2 Analysis
Why Dynamic Layer Selection. To analyze the impact of intermediate layers, we train UGround on the ReasonSeg train set with two strategies: (1) relying solely on the last hidden layer ($\ell=L$), and (2) dynamically selecting from the intermediate layers ($\ell\in\{1,\dots,L-1\}$). Qualitatively, we visualize the similarity maps across layers 1–40 of $\mathcal{G}_{\mathcal{T}}$ in Fig. 4. We observe that the fixed-last-layer strategy produces noisy similarity maps in intermediate layers (Fig. 4(a)), whereas the intermediate-layer counterpart yields a more discriminative foreground (Fig. 4(b)). Quantitatively, we plot the predicted cIoU of the embeddings across layers 1–40. Fig.2(a) shows that every layer from 10–40 improves over the fixed-last-layer strategy by a clear and consistent margin. We further plot the per-layer similarity-map loss with respect to the ground-truth mask. Fig.2(b) shows that the intermediate-layer strategy begins converging from layer 19 and ultimately reaches a lower loss level, while the fixed-last-layer strategy only starts converging from layer 28, indicating that dynamic layer selection accelerates convergence. Dynamic layer selection provides insight into the behavior of intermediate layers. By allowing these layers to directly participate in interactions and joint training with SAM, akin to “cheating” in a phone game (i.e., receiving intermediate hints), each layer’s performance is improved, ultimately resulting in an overall enhancement.
Figure 2: (a) shows that the dynamic layer selection strategy outperforms the fixed last hidden layer strategy across all middle layers. (b) plots the loss of similarity maps against the soft ground-truth mask, demonstrating faster convergence at intermediate layers.
Why Similarity as Mask. To investigate whether the similarity map can serve as a prompt, we conduct two analyses. First, to measure the consistency ($\mathcal{M}_{\mathrm{IoU}}$) between the similarity map and the ground-truth mask, we compute the grid search-based Intersection over Union (IoU) following READ (Qian et al., 2025). Second, to probe whether SAM can understand the semantics encoded in the similarity map, we directly use the similarity map as a logit mask to prompt the original SAM (Kirillov et al., 2023) (without any additional training), denoted by $\mathcal{M}_{\mathrm{prompt}}$. Table 2 shows that SAM can potentially leverage the similarity map as a prompt ($17.0\%$ vs. $48.4\%$). Further, when the similarity map is directly converted into a binary mask for prediction, its performance measured by $\mathcal{M}_{\mathrm{IoU}}$ can even rival that of the trained <SEG>prompt, as evidenced by a 4.3% higher cIoU on the ReasonSeg test set ($35.0\%\ \text{vs.}\ 30.7\%$) in the $3^{\text{rd}}$ row.
Table 2: Quantitative analysis of the similarity maps on the ReasonSeg test split. $\mathcal{M}$prompt indicates mask as a prompt for the original SAM, while $\mathcal{M}$IoU quantifies the overlap between the similarity map and the ground-truth mask. <SEG>prompt refers to <SEG> token as prompt for SAM.
Method $\mathcal{M}$prompt $\mathcal{M}$IoU <SEG>prompt
@@
Summary. Analysis of the similarity maps reveals the following: (a) Similarity maps in LMMs are highly consistent with the ground-truth mask, suggesting that they can potentially serve as prompts for generating the segmentation mask. (b) Dynamic layer selection allows the intermediate layers to directly interact with SAM, resulting in a more discriminative representation, as reflected by the activated regions in the similarity maps. Motivated by the insights from the similarity maps, we leverage it in three ways: as a logit mask to prompt SAM (prompt), as a loss constraint to guide the model on where to “attend” (constraint), and as a policy reward for dynamic layer selection (signal) (see Appendix D for further analysis).
@@ 4.1 Policy-Prompted Masking
$\pi_{\theta}(\ell\mid\mathcal{H}_{t^{*}})=\frac{\exp(s_{\ell})}{\sum_{j=1}^{L}\exp(s_{j})},s_{\ell}=\bm{h}_{t^{*}}^{(\ell)}\cdot\mathbf{w}_{\ell},\vskip-8.5359pt$ (4)
@@
where the log-probability of the sampled action $a\equiv\ell^{*}$ is weighted by the advantage rbr-b. To get the reward $r$, we first compute the similarity map $\mathcal{M}$ between the <SEG> token and image tokens at layer $\ell$ and its corresponding soft mask $\mathcal{M}_{\sigma}$. Consider a sequence of discrete image tokens $\mathbf{z}=(z_{1},z_{2},\dots,z_{k})$ in the input sequence $\mathbf{x}$, where $k$ is the number of image tokens. The image token embeddings across layers are then denoted as
$\mathcal{H}_{\mathbf{z}}=\{\bm{h}_{z}^{(\ell)}\mid z\in\mathbf{z},\ \ell=1,\dots,L\}\in\mathbb{R}^{k\times L\times d}$, where $\bm{h}_{z}^{(\ell)}$ is the embedding of the $z$-th image token at layer $\ell$. Given the <SEG> token embedding $\bm{h}_{t^{*}}^{(\ell)}\in\mathbb{R}^{d}$ and the image token embeddings $\mathcal{H}_{\mathbf{z}}^{(\ell)}\in\mathbb{R}^{k\times d}$ at the selected layer $\ell$, the similarity scores can be computed as
@@
where each element $\mathcal{S}_{i}^{(\ell)}=(h_{z_{i}}^{(\ell)})^{\top}h_{t^{*}}^{(\ell)}$ represents the similarity between the <SEG> token and the $i$-th selected image token at layer $\ell$. $\mathcal{S}^{(\ell)}$ are arranged in a 2D grid, and interpolated to obtain the soft logit mask $\mathcal{M}\in[0,1]^{H\times W}$. Let the ground-truth mask be $M\in\{0,1\}^{H\times W}$, which we regard as a function $M:\{1,\dots,H\}\times\{1,\dots,W\}\to\{0,1\}$ over the discrete pixel grid. We extend $M$ to the continuous domain $\mathbb{R}^{2}$ by assigning each pixel a constant value on its corresponding region in the plane. Applying Gaussian smoothing to this extended function yields a continuous-valued function $M_{\sigma}:\mathbb{R}^{2}\to[0,1]$, where $M_{\sigma}$ serves as the soft label (heatmap). Formally, $M_{\sigma}$ can be expressed as a convolution integral as $M_{\sigma}(u,v)=\iint_{\mathbb{R}^{2}}M(x,y)\,G_{\sigma}(u-x,v-y)\,dx\,dy,$ where $G_{\sigma}(x,y)\propto\exp(-\frac{x^{2}+y^{2}}{2\sigma^{2}})$, is a 2D Gaussian kernel with standard deviation $\sigma$. The reward for REINFORCE is computed using the similarity map $\mathcal{M}$ and the soft label $M_{\sigma}$ as $r=-\big(\mathcal{L}_{bce}(\mathcal{M},M_{\sigma})+\mathcal{L}_{dice}(\mathcal{M},M_{\sigma})\big),$ where $\mathcal{L}_{bce}$ and $\mathcal{L}_{dice}$ denote the Gaussian BCE and Dice losses defined over the similarity map $\mathcal{M}$. The policy loss is then defined as